HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

AB INITIO LINE LIST FOR DISULFUR MONOXIDE UP TO 1800 CM^{−1}

OLEG EGOROV, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia, Faculty of Physics Tomsk State University 36, Lenin Ave., Tomsk, 634050 Russia; \underline{MICHAËL REY} , Groupe de Spectrométrie Moléculaire et Atmosphérique UMR CNRS 7331, UFR Sciences BP 1039, 51687 Reims Cedex 2, France; ROMAN V. KOCHANOV, ANDREI V. NIKITIN, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk,

634055 Russia, Tomsk State University 36, Lenin Ave., Tomsk, 634050 Russia; \underline{VLADIMIR TYUTEREV} , Tomsk State University 36, Lenin Ave., Tomsk, 634050 Russia, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric

Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia

The line list for the main isotopologue of disulfur monoxide (^{32}S_{2}^{16}O) was variationally computed using both accurate ab initio potential energy (PES) and dipole moment surfaces (DMS). The applied ab initio PES^{1} allows to describe the experimental band origins of^{32}S_{2}^{16(18)}O within the error of 0.5 cm^{−1}. The ab initio DMS was constructed via the finite difference approximation using the CCSD(T)/aug-cc-pV5Z approach. The final line list covers the 0-1800 cm^{−1} region and contains more than one million transitions with the intensity cutoff value of 10^{−25} cm/molecule. The simulated absorbances agree well with the available measurements, particularly, in the regions of the ν_{2} (380 cm^{−1})^{2} and ν_{3} (679 cm^{−1})^{3} fundamental bands. The integrated intensities of the strongest cold and hot bands located in the 0-1800 cm^{−1} region are presented.

This work was supported by a Grant from the RF Ministry of Education and Science for the project SACHA: (N 075-15-2021-1412, contract identifier RF2251.62321X0012) in a framework of bilateral French-Russian “Kolmogorov” (PHC) cooperation program (2021-2251-ΠΠ408) for Science and Technology. The support from the ROMEO computer center of Reims-Champagne-Ardenne as well as from the French-Russian collaboration program IRP SAMIA2 is also acknowledged.

1doi:10.1016/j.cplett.2022.140216, O. Egorov, M. Rey, R. V. Kochanov, A. V. Nikitin, V. Tyuterev, Chem. Phys. Lett., 811, 140216, (2023).

2doi:10.1016/j.jms.2015.02.014, M.A. Martin-Drumel, C.P. Endres, O. Zingsheim et al. J. Mol.

Spectrosc., 315, 72-79, (2015).

3doi:10.1016/j.jms.2015.12.009, S. Thorwirth, M.A. Martin-Drumel, C.P. Endres et al. J. Mol.

Spectrosc., 319, 47-49, (2016).

p-number: p102  Submitted on Tue, 13 Jun 2023 15:37:53 +0200